OpenStudy (anonymous):

HELP PLEASE !!(: Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. A) y = x + 8 B) y = x - 4 C) y = x + 4 D) y = x - 8

OpenStudy (anonymous):

m=1, y+2= x-6 y=x-8

OpenStudy (anonymous):

thank you so soo much